# Polynomial-Time Preparation of Low-Temperature Gibbs States for 2D Toric Code

### v1

@@ 1.4 Organization
In the following part of the paper, we start with a brief introduction to the Davies generator and properties of its spectral gap in Section 2 . A technical overview of the proof of 1 is provided in Section 3. The detailed introduction to the 2D toric code and its proof can be found in Section 4. Additionally, in Appendix A, we discuss a simpler case of the 1D ferromagnetic Ising chain for completeness.
@@ Acknowledgment
This material is partially supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (Z.D.), by the Challenge Institute for Quantum Computation (CIQC) funded by National Science Foundation (NSF) through grant number OMA-2016245 (L.L.), and by DOE Grant No. DE-SC0024124 (R.Z.). L.L. is a Simons Investigator in Mathematics. We thank Garnet Chan, Li Gao, Yunchao Liu, John Preskill for helpful discussions and feedbacks.
@@ 2 Preliminaries
where the transition rate function $\gamma_{a}(\omega)>0$ is given by the Fourier transform of the bath auto-correlation function satisfying the KMS condition [32]:
@@
It follows that $\mathcal{L}_{\beta}$ is similar to a self-adjoint operator for the HS inner product, called the master Hamiltonian, and has only real spectrum. To be precise, we introduce the transform $\varphi_{X}:=X{\sigma_{\beta}}^{1/2}$, which gives $\langle Y,X\rangle_{\sigma_{\beta}}=\langle\varphi_{Y},\varphi_{X}\rangle$. Then, the master Hamiltonian ~$\widetilde{\mathcal{L}}_{\beta}$ is given by the similar transform of $\mathcal{L}_{\beta}$ via $\varphi_{X}$:
@@
${\rm Gap}(\mathcal{L}_{\beta})=\inf_{X\neq 0\,,\operatorname{Tr}(\sigma_{\beta}X)=0}\frac{\langle X,\mathcal{L}_{\beta}(X)\rangle_{\sigma_{\beta}}}{\langle X,X\rangle_{\sigma_{\beta}}}\,.$
@@
$\left\lVert\mathcal{L}_{\beta}\right\rVert_{\sigma_{\beta}\to\sigma_{\beta}}=\sup_{X\neq 0}\frac{\langle X,\mathcal{L}_{\beta}(X)\rangle_{\sigma_{\beta}}}{\langle X,X\rangle_{\sigma_{\beta}}}\,.$
By [54, 57], a sufficient and necessary condition for the primitivity of $\mathcal{L}_{\beta}$ is the $\mathbb{C}$-algebra generated by all the jump operators $\{S_{a}(\omega)\}_{a,w}$ is the whole algebra $\mathcal{B}(\mathcal{H})$. By this condition, one can check that for the choice of $\{S_{a}\}_{a\in\mathcal{A}}=\{\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}_{i=1}^{N}$, the associated Davies generator is always primitive. Without loss of generality, when discussing the Gibbs samplers in Appendix A and Section 4, we always let $\{\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}_{i=1}^{N}$ be a subset of $\{S_{a}\}_{a\in\mathcal{A}}$ to guarantee the primitivity.
@@
The second statement in Lemma 3 means that for any primitive Davies generator with $\sigma_{\beta}$ being invariant, adding any other Davies generator keeping the invariant state can only increase the spectral gap.
3 Technical overview
  1. (a)
  2. (b)
    Demonstrate efficient transition between logic subspaces.
  3. (c)
    Demonstrate fast mixing inside the syndrome subspace.
Thus, for a lower bound estimate of $\mathrm{Gap}(-\mathcal{L}^{\rm gapped})$, it suffices to consider $\mathrm{Gap}(-\mathcal{L}^{\rm gapped}|_{\mathcal{B}_{{\bf 1},{\bf 1}}})$ and $\lambda_{\min}(-\mathcal{L}^{\rm gapped})|_{\mathcal{B}_{B_{1},B_{2}}}$ with $B_{1}\neq{\bf 1}\ \text{or}\ B_{2}\neq{\bf 1}$. The latter term characterizes the transition rate between different logical subspaces. If $\mathcal{L}_{\beta}$ contains only local coupling operators, the system requires a long evolution time to overcome the energy barrier and transit between different logical subspaces, which leads to a slow mixing time scaling as $\exp(\Theta(\beta))$ according to [2]. In our work, an important observation is that the global logical operators $\mathsf{X}_{1},\mathsf{Z}_{1},\mathsf{X}_{2},\mathsf{Z}_{2}$ can directly flip the logical qubits, enabling transitions between different logical subspaces without the need to overcome the energy barrier. The efficient transition in logic density space implies a fast decaying of $\exp(t\mathcal{L}^{\rm gapped})$ in the logic subspace $\{\mathcal{B}_{B_{1},B_{2}}\}_{B_{1}\neq{\bf 1}\ \text{or}\ B_{2}\neq{\bf 1}}$ and a lower bound of $\{-\mathcal{L}^{\rm gapped}|_{\mathcal{B}_{B_{1},B_{2}}}\}_{B_{1}\neq{\bf 1}\ \text{or}\ B_{2}\neq{\bf 1}}$. Specifically, in 8 in Section 4.2, we show
According to the above analysis, the remaining thing to lower bound the spectral gap of $\mathcal{L}_{\beta}$ is to study the spectral gap of $\mathcal{L}_{\rm local}$ on $\mathcal{B}_{{\bf 1},{\bf 1}}$. Roughly speaking, this requires us to prove 2. Similar task is done in [2]. However, we emphasize that the lower bound and proof technique in [2] is not suitable for our purpose. In [2, Proposition 2], while the spectral gap is independent of the system size, it decays as $\exp(-\Theta(\beta))$, indicating slow mixing in the syndrome subspace at low temperatures. A main contribution of this work is to show that this lower bound is not tight when $\beta\gg 1$, and the system actually mixes fast in the syndrome subspace at low temperatures. To this end, we develop a new iteration argument and a decomposition trick to prove that $\mathrm{Gap}\left(\mathcal{L}_{\rm local}\middle|_{\mathcal{B}_{{\bf 1},{\bf 1}}}\right)\geq\min\left\{\exp(-\Theta(\beta)),\mathrm{poly}(1/N)\right\}$, which provides a much sharper lower bound of the spectral gap in the lower temperature regime. This result is summarized in 9 in Section 4.2.1, which also provides a proof of 2. Plugging this into (3.1) and (3.2), we can conclude 1.
@@ 3 Technical overview
Figure 1: 2D toric code. Each blue plaquette contains one 4-local operator $\textbf{Z}_{p}=\prod_{i\,\in\,\text{plaquette }p}\sigma^{z}_{i}$. Each red star contains one 4-local operator $\textbf{X}_{s}=\prod_{i\,\in\,\text{star }s}\sigma^{x}_{i}$. The dashed line on the right/bottom is the same as the solid line on the left/up to indicate the periodic boundary condition.
@@ 4.1 Ground states and observable algebra
Analogous to the 1D ferromagnetic Ising chain (see Appendix A), the 2D toric code Hamiltonian $H^{\text{toric}}$ is frustration-free, i.e., its ground state is simultaneously an eigenvector with eigenvalue 1 for all local terms $\mathbf{X}_{s}$ and $\mathbf{Z}_{p}$. Due to the toric structure with periodic boundary conditions, there are two topologically protected degrees of freedom (i.e., two logical qubits), resulting in a four-dimensional ground state space.
@@
$\mathsf{X}_{1}:=\prod_{j\in\text{orange dots}}\sigma^{x}_{j}\,,\quad\mathsf{Z}_{1}:=\prod_{j\in\text{green squares}}\sigma^{z}_{j}\,,\quad\mathsf{Y}_{1}:=i\mathsf{Z}_{1}\mathsf{X}_{1}\,,$ (4.4)
$\mathsf{X}_{2}:=\prod_{j\in\text{blue dots}}\sigma^{x}_{j}\,,\quad\mathsf{Z}_{2}:=\prod_{j\in\text{red squares}}\sigma^{z}_{j}\,,\quad\mathsf{Y}_{2}:=i\mathsf{Z}_{2}\mathsf{X}_{2}\,.$
Here, $\{\mathsf{X}_{1},\mathsf{Y}_{1},\mathsf{Z}_{1}\}$ and $\{\mathsf{X}_{2},\mathsf{Y}_{2},\mathsf{Z}_{2}\}$ generates two commuting Pauli algebras. Moreover, the operators $\mathsf{X}_{1}$ and $\mathsf{X}_{2}$ give one-to-one correspondences between four topological equivalent classes [3, Fig. 3.2] (see also [12, Section 3.2]). Specifically, we consider four orthogonal vectors that satisfy $\textbf{Z}_{p}\ket{\phi}=\ket{\phi}$:
@@
    $\textbf{Z}_{p_{1}/p_{2}}\ket{\psi_{p_{1},p_{2}}}=-\ket{\psi_{p_{1},p_{2}}},\quad H^{\rm toric}\ket{\psi_{p_{1},p_{2}}}=E_{\rm ground}+2\,,$
@@
    $\textbf{X}_{s_{1}/s_{2}}\ket{\psi_{s_{1},s_{2}}}=-\ket{\psi_{s_{1},s_{2}}},\quad H^{\rm toric}\ket{\psi_{s_{1},s_{2}}}=E_{\rm ground}+2\,.$
    thanks to $\textbf{X}_{s_{1}/s_{2}}W_{l}^{e}=-W_{l}^{e}\textbf{X}_{s_{1}/s_{2}}$.
@@
These elementary excited states |$\ket{\psi_{p_{1},p_{2}}},\ket{\psi_{s_{1},s_{2}}}$ are often called quasi-particles pairs (or quasi-particles for short). All excited states of $H^{\rm toric}$ can be expressed using these quasi-particles.
@@
associated with the decomposition (4.8), where the algebras $\mathcal{A}^{\rm full}_{\rm m}$ and $\mathcal{A}^{\rm full}_{\rm e}$ are generated by $\{\mathbf{Z}_{p}\}_{p}\cup\{\sigma^{x}_{j}\}_{j\in\text{\rm snake}}$ and $\{\mathbf{X}_{s}\}_{s}\cup\{\sigma^{z}_{j}\}_{j\in\text{\rm comb}}$. In addition, the four subalgebras $\mathcal{Q}_{1}$, $\mathcal{Q}_{2}$, $\mathcal{A}_{\rm m}^{\rm full}$, and $\mathcal{A}_{\rm e}^{\rm full}$ are commutative with each other.
To give further interpretation on the decomposition (4.11), we define four subspaces $\mathcal{H}_{g}$ with $g=o,\ |,\ -,\ +$ spanned by the ground state |$\ket{\psi_{g}}$ in (4.5) and its excited states $W_{l}^{m/e}\ket{\psi_{g}}$ via (4.6) and (4.7). Then it is easy to see $\dim(\mathcal{H}_{g})=2^{N-2}$ with $N=2L^{2}$. The algebra $\mathcal{Q}_{1}\otimes\mathcal{Q}_{2}$ gives all the linear transformation between subspaces $\mathcal{H}_{g}$ with $g=o,\ |,\ -,\ +$, while on each subspace $\mathcal{H}_{g}$, the linear maps are characterized by $\mathcal{A}^{\rm full}_{\rm m}\otimes\mathcal{A}^{\rm full}_{\rm e}$. In particular, recalling the basis (4.10), $\mathcal{A}^{\rm full}_{\rm m}$ (resp., $\mathcal{A}^{\rm full}_{\rm e}$) transfers the bond |$\ket{m}$ (resp., |$\ket{e}$), for example,
@@ 4.2 Spectral gap of the Gibbs sampler
Here $\mathcal{L}^{\rm rest}:=\mathcal{L}_{\text{\rm local full}}-\mathcal{L}_{\rm local}$ is a local Lindbladian with other Pauli couplings not included in $\mathcal{L}_{\rm local}$. From Lemma 3 (item 2), it suffices to limit our discussion to the part $\mathcal{L}^{\rm gapped}$ and show that it is primitive with the desired spectral gap lower bound $\max\left\{\exp(-\mathcal{O}(\beta)),\Omega(N^{-3})\right\}$.
@@
due to $P_{j}^{\pm}\sigma_{j}^{x}P_{j}^{0}=P^{+}_{j}\sigma_{j}^{x}P^{+}_{j}=P^{-}_{j}\sigma_{j}^{x}P^{-}_{j}=0$, where $a_{j}$, $a_{j}^{\dagger}$, $a_{j}^{0}$ correspond to the Bohr frequencies $-4$, $4$, 00 of the Hamiltonian (4.1), respectively. Then, we can write $\mathcal{L}_{\sigma_{j}^{x}}$ by Eqs. 2.3 and 2.7 with Glauber-type transition (2.5):
@@
since $\mathsf{O}$ has only the component with Bohr frequency zero. It is clear that
$\left\lVert\mathcal{L}_{\mathsf{O}}\right\rVert_{\sigma_{\beta}\to\sigma_{\beta}}=\Theta(1)\,,\quad\text{uniformly in $\beta$}\,.$ (4.20)

### v2

@@ 1.4 Organization
In the following part of the paper, we start with a technical overview of the proof of 1 in Section 2. Then, we provide a brief introduction to the Davies generator and properties of its spectral gap in Section 3. The detailed introduction to the 2D toric code and its proof can be found in Section 4. Additionally, in Appendix B, we discuss a simpler case of the 1D ferromagnetic Ising chain for completeness.
@@ Acknowledgment
This work is supported by the National Key R&\&&D Program of China Grant No. 2024YFA1016000 (B.L.), the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (Z.D.), by the Challenge Institute for Quantum Computation (CIQC) funded by National Science Foundation (NSF) through grant number OMA-2016245 (L.L.), and by DOE Grant No. DE-SC0024124 (R.Z.). L.L. is a Simons Investigator in Mathematics. We thank Garnet Chan, Li Gao, Jiaqing Jiang, Yunchao Liu, and John Preskill for helpful discussions and feedback.
2 Technical overview
  1. (a)
  2. (b)
    Demonstrate efficient transition between logic subspaces.
  3. (c)
    Demonstrate fast mixing inside the syndrome subspace.
Thus, for a lower bound estimate of $\mathrm{Gap}(-\mathcal{L}^{\rm gapped})$, it suffices to consider $\mathrm{Gap}(-\mathcal{L}^{\rm gapped}|_{\mathcal{B}_{{\bf 1},{\bf 1}}})$ and $\lambda_{\min}(-\mathcal{L}^{\rm gapped})|_{\mathcal{B}_{B_{1},B_{2}}}$ with $B_{1}\neq{\bf 1}\ \text{or}\ B_{2}\neq{\bf 1}$. The latter term characterizes the transition rate between different logical subspaces. If $\mathcal{L}_{\beta}$ contains only local coupling operators, the system requires a long evolution time to overcome the energy barrier and to transit between different logical subspaces, which leads to a slow mixing time scaling as $\exp(\Theta(\beta))$ according to [2]. In our work, an important observation is that the global logical operators $\mathsf{X}_{1},\mathsf{Z}_{1},\mathsf{X}_{2},\mathsf{Z}_{2}$ can directly flip the logical qubits, enabling transitions between different logical subspaces without the need to overcome the energy barrier. The efficient transition in logic density space implies a fast decaying of $\exp(t\mathcal{L}^{\rm gapped})$ in the logic subspace $\{\mathcal{B}_{B_{1},B_{2}}\}_{B_{1}\neq{\bf 1}\ \text{or}\ B_{2}\neq{\bf 1}}$ and a lower bound of $\{-\mathcal{L}^{\rm gapped}|_{\mathcal{B}_{B_{1},B_{2}}}\}_{B_{1}\neq{\bf 1}\ \text{or}\ B_{2}\neq{\bf 1}}$. Specifically, in 8 in Section 4.2, we show
According to the above analysis, the remaining thing to lower bound the spectral gap of $\mathcal{L}_{\beta}$ is to study the spectral gap of $\mathcal{L}_{\rm local}$ on $\mathcal{B}_{{\bf 1},{\bf 1}}$. Roughly speaking, this requires us to prove 2. A similar task is done in [2]. However, we emphasize that the lower bound and proof technique in [2] is not suitable for our purpose. In [2, Proposition 2], while the spectral gap is independent of the system size, it decays as $\exp(-\Theta(\beta))$, indicating slow mixing in the syndrome subspace at low temperatures. A main contribution of this work is to show that this lower bound is not tight when $\beta\gg 1$, and the system actually mixes fast in the syndrome subspace at low temperatures. To this end, we develop a new iteration argument, a decomposition trick, and a tight estimation of the minimal eigenvalue of a perturbed graph Laplacian to prove that $\mathrm{Gap}\left(\mathcal{L}_{\rm local}\middle|_{\mathcal{B}_{{\bf 1},{\bf 1}}}\right)\geq\min\left\{\exp(-\Theta(\beta)),\mathrm{poly}(1/N)\right\}$, which provides a much sharper lower bound of the spectral gap in the lower temperature regime. This result is summarized in 9 in Section 4.2.1, which also provides a proof of 2. Plugging this into (2.1) and (2.2), we can conclude 1.
@@ 3 Preliminaries
where the transition rate function $\gamma_{a}(\omega)>0$ is given by the Fourier transform of the bath autocorrelation function satisfying the KMS condition [35]:
@@
It follows that $\mathcal{L}_{\beta}$ is similar to a self-adjoint operator for the HS inner product, called the master Hamiltonian, and has only the real spectrum. To be precise, we introduce the transform $\varphi_{X}:=X{\sigma_{\beta}}^{1/2}$, which gives $\langle Y,X\rangle_{\sigma_{\beta}}=\langle\varphi_{Y},\varphi_{X}\rangle$. Then, the master Hamiltonian ~$\widetilde{\mathcal{L}}_{\beta}$ is given by the similar transform of $\mathcal{L}_{\beta}$ via $\varphi_{X}$:
@@
${\rm Gap}(\mathcal{L}_{\beta})=\inf_{X\neq 0\,,\operatorname{Tr}(\sigma_{\beta}X)=0}\frac{\langle X,-\mathcal{L}_{\beta}(X)\rangle_{\sigma_{\beta}}}{\langle X,X\rangle_{\sigma_{\beta}}}\,.$
@@
$\left\lVert\mathcal{L}_{\beta}\right\rVert_{\sigma_{\beta}\to\sigma_{\beta}}=\sup_{X\neq 0}\frac{\langle X,-\mathcal{L}_{\beta}(X)\rangle_{\sigma_{\beta}}}{\langle X,X\rangle_{\sigma_{\beta}}}\,.$ (3.10)
By [57, 60], a sufficient and necessary condition for the primitivity of $\mathcal{L}_{\beta}$ is the $\mathbb{C}$-algebra generated by all the jump operators $\{S_{a}(\omega)\}_{a,w}$ is the whole algebra $\mathcal{B}(\mathcal{H})$. By this condition, one can check that for the choice of $\{S_{a}\}_{a\in\mathcal{A}}=\{\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}_{i=1}^{N}$, the associated Davies generator is always primitive. Without loss of generality, when discussing the Gibbs samplers in Appendix B and Section 4, we always let $\{\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}_{i=1}^{N}$ be a subset of $\{S_{a}\}_{a\in\mathcal{A}}$ to guarantee the primitivity.
@@
The second statement in Lemma 3 means that for any primitive Davies generator with $\sigma_{\beta}$ being invariant, adding any other Davies generator keeping the invariant state does not decrease the spectral gap.
@@
Figure 1: 2D toric code. Each blue plaquette contains one 4-local operator $\textbf{Z}_{p}=\prod_{i\,\in\,\text{plaquette }p}\sigma^{z}_{i}$. Each red star contains one 4-local operator $\textbf{X}_{s}=\prod_{i\,\in\,\text{star }s}\sigma^{x}_{i}$. The dashed line on the right/bottom is the same as the solid line on the left/top to indicate the periodic boundary condition.
@@ 4.1 Ground states and observable algebra
Analogous to the 1D ferromagnetic Ising chain (see Appendix B), the 2D toric code Hamiltonian $H^{\text{toric}}$ is frustration-free, i.e., its ground state is simultaneously an eigenvector with eigenvalue 1 for all local terms $\mathbf{X}_{s}$ and $\mathbf{Z}_{p}$. Due to the toric structure with periodic boundary conditions, there are two topologically protected degrees of freedom (i.e., two logical qubits), resulting in a four-dimensional ground state space.
@@
$\mathsf{X}_{1}:=\prod_{j\in\text{orange dots}}\sigma^{x}_{j}\,,\quad\mathsf{Z}_{1}:=\prod_{j\in\text{green squares}}\sigma^{z}_{j}\,,\quad\mathsf{Y}_{1}:={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-i}\mathsf{Z}_{1}\mathsf{X}_{1}\,,$ (4.4)
$\mathsf{X}_{2}:=\prod_{j\in\text{blue dots}}\sigma^{x}_{j}\,,\quad\mathsf{Z}_{2}:=\prod_{j\in\text{red squares}}\sigma^{z}_{j}\,,\quad\mathsf{Y}_{2}:={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-i}\mathsf{Z}_{2}\mathsf{X}_{2}\,.$
Here, $\{\mathsf{X}_{1},\mathsf{Y}_{1},\mathsf{Z}_{1}\}$ and $\{\mathsf{X}_{2},\mathsf{Y}_{2},\mathsf{Z}_{2}\}$ generates two commuting Pauli algebras. Moreover, the operators $\mathsf{X}_{1}$ and $\mathsf{X}_{2}$ give one-to-one correspondences between four topological equivalent classes [3, Fig. 3.2] (see also [13, Section 3.2]). Specifically, we consider four orthogonal vectors that satisfy $\textbf{Z}_{p}\ket{\phi}=\ket{\phi}$:
@@
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}_{p_{1}}\ket{\psi_{p_{1},p_{2}}}=\textbf{Z}_{p_{2}}\ket{\psi_{p_{1},p_{2}}}=-\ket{\psi_{p_{1},p_{2}}},\quad H^{\rm toric}\ket{\psi_{p_{1},p_{2}}}=(E_{\rm ground}+2)\ket{\psi_{p_{1},p_{2}}}\,,}$
@@
    $\textbf{X}_{s_{1}}\ket{\psi_{s_{1},s_{2}}}=\textbf{X}_{s_{2}}\ket{\psi_{s_{1},s_{2}}}=-\ket{\psi_{s_{1},s_{2}}},\quad H^{\rm toric}\ket{\psi_{s_{1},s_{2}}}=(E_{\rm ground}+2)\ket{\psi_{s_{1},s_{2}}}\,.$
    thanks to $\textbf{X}_{s_{1}(\text{or}\ s_{2})}W_{l}^{e}=-W_{l}^{e}\textbf{X}_{s_{1}(\text{or}\ s_{2})}$.
@@
These elementary excited states |$\ket{\psi_{p_{1},p_{2}}},\ket{\psi_{s_{1},s_{2}}}$ are often called quasi-particle pairs (or quasi-particles for short). All excited states of $H^{\rm toric}$ can be expressed using these quasi-particles.
@@
associated with the decomposition (4.8), where the algebras $\mathcal{A}^{\rm full}_{\rm m}$ and $\mathcal{A}^{\rm full}_{\rm e}$ are generated by $\{\mathbf{Z}_{p}\}_{p}\cup\{\sigma^{x}_{j}\}_{j\in\text{\rm snake}}$ and $\{\mathbf{X}_{s}\}_{s}\cup\{\sigma^{z}_{j}\}_{j\in\text{\rm comb}}$. In addition, the four subalgebras $\mathcal{Q}_{1}$, $\mathcal{Q}_{2}$, $\mathcal{A}_{\rm m}^{\rm full}$, and $\mathcal{A}_{\rm e}^{\rm full}$ are commutative with each other, namely, the operators in these subalgebras acting on the entire Hilbert space $\mathcal{H}$ mutually commute with each other.
To give further interpretation on the decomposition (4.11), we define four subspaces $\mathcal{H}_{g}$ with $g=o,\ |,\ -,\ +$ spanned by the ground state |$\ket{\psi_{g}}$ in (4.5) and its excited states $W_{l}^{m/e}\ket{\psi_{g}}$ via (4.6) and (4.7). Then it is easy to see $\dim(\mathcal{H}_{g})=2^{N-2}$ with $N=2L^{2}$. The algebra $\mathcal{Q}_{1}\otimes\mathcal{Q}_{2}$ gives all the linear transformation between subspaces $\mathcal{H}_{g}$ with $g=o,\ |,\ -,\ +$, while on each subspace $\mathcal{H}_{g}$, the linear maps are characterized by $\mathcal{A}^{\rm full}_{\rm m}\otimes\mathcal{A}^{\rm full}_{\rm e}$. In particular, recalling the basis (4.10), $\mathcal{A}^{\rm full}_{\rm m}$ (resp., $\mathcal{A}^{\rm full}_{\rm e}$) acts only on |$\ket{m}$ (resp., |$\ket{e}$), for example,
@@ 4.2 Spectral gap of the Gibbs sampler
Here $\mathcal{L}^{\rm rest}:=\mathcal{L}_{\text{\rm local full}}-\mathcal{L}_{\rm local}$ is a local Lindbladian with other Pauli couplings not included in $\mathcal{L}_{\rm local}$. From Lemma 3 (item 2), it suffices to limit our discussion to the part $\mathcal{L}^{\rm gapped}$ and show that it is primitive with the desired spectral gap lower bound $\max\left\{\exp(-\mathcal{O}(\beta)),\Omega(N^{-2})\right\}$.
@@
due to $P_{j}^{\pm}\sigma_{j}^{x}P_{j}^{0}=P^{+}_{j}\sigma_{j}^{x}P^{+}_{j}=P^{-}_{j}\sigma_{j}^{x}P^{-}_{j}=0$, where $a_{j}$, $a_{j}^{\dagger}$, $a_{j}^{0}$ correspond to the Bohr frequencies $-4$, $4$, 00 of the Hamiltonian (4.1), respectively. Then, we can write $\mathcal{L}_{\sigma_{j}^{x}}$ by Eqs. 3.3 and 3.7 with Glauber-type transition (3.5):
@@
since $\mathsf{O}$ has only the component with Bohr frequency zero. According to (3.10) and (4.19), it is straightforward to see that
$\left\lVert\mathcal{L}_{\mathsf{O}}\right\rVert_{\sigma_{\beta}\to\sigma_{\beta}}=\Theta(\|\mathsf{O}\|_{\rm op}^{2})=\Theta(1)\,,\quad\text{uniformly in $\beta$}\,,$ (4.20)
where $\left\lVert\cdot\right\rVert_{\rm op}$ denotes the standard operator norm.